# REArtGS: Reconstructing and Generating Articulated Objects via 3D Gaussian Splatting with Geometric and Motion Constraints

### v3

@@ 4.1 Experimental Setting
Datasets. We conduct surface reconstruction experiments on PartNet-Mobility [31] and AKB-48 [16] to evaluate the reconstruction quality of our method. The PartNet-Mobility dataset is a synthetic dataset comprising articulated objects across 46 categories, characterized by a diverse scale and geometry. Following PARIS, we selected the same 10 categories for our experiments. Each category comprises 64 to 100 RGB views at two arbitrary states, which are sampled randomly from their upper hemispheres. AKB-48 is a real-world dataset containing an extensive collection of articulated objects scanned in real indoor scenes. We select 5 categories with complicated geometry and texture and sample 100 RGB views as training inputs.
Metrics. Chamfer Distance (CD), F1-score and Earth Mover’s Distance (EMD) [40] are usually adopted as evaluation metrics for surface quality. However, these methods typically focus on the whole surface, endowing the unseen areas (such as the bottom of the object) with considerable weights during evaluation. This is unreasonable since we tend to be more concerned with the upper hemisphere regions of articulated objects. Therefore, we sample point clouds from the fragments between the rays cast by test cameras and the surface mesh to calculate EMD and Chamfer-L1 distance, defined as CD (rs). We also evaluate the F1-score and Chamfer-L1 distance for the whole surface, defined as CD (ws), to comprehensively evaluate the surface quality. Note that the CD values are multiplied by 1000 and the distance threshold of F1-score is set to 0.4. The calculation of these metrics can be referred to in our supplementary materials.
All the experiments are conducted on a single RTX 4090 GPU. We also provide detailed implementation of our approach and baselines in the supplementary materials.
@@ 4.2 Mesh Reconstruction Performance
We report the quantitative results of mesh reconstruction in Table. 1. As can be observed, our REArtGS outperforms state-of-the-art approaches in the mean across all metrics with 3.79, 1.236, 0.294 and 0.695 respectively. Compared to A-SDF [21] and Ditto [10] that exploits 3D inputs for articulated object surface reconstruction, our method achieves higher reconstruction quality in the majority of categories. Although ditto surpasses our approach in several categories (Washer, Oven, and Storage), it employs additional 3D data for pre-training and requires the united robotics description format (URDF) files of articulated objects. However, our method merely requires the RGB view collections at two arbitrary states, which is more convenient in practical applications. In terms of PARIS that utilizes the neural implicit representation for surface reconstruction, our approach exhibits significantly smoother and clearer surfaces, as shown in Fig. 4. Although PARIS slightly outperforms our approach in a few categories of CD (ws), it exhibits considerable disadvantages compared with our approach in CD (rs) and EMD scores, which better evaluate the substantial surface quality. As for the 3DGS-based reconstruction method GOF, our approach achieves marked superiority in the vast majority of categories. This is primarily attributed to our geometric constraints regularizing the opacity fields and facilitating more reasonable distribution of 3D Gaussian primitives over scene surfaces.
@@ 4.3 Mesh Generation Performance
We report the quantitative results of mesh reconstruction in Table. 2 and present the qualitative results Fig. 5. Please refer to our supplementary material for more qualitative results. Our REArtGS achieves the best mean results across all metrics with 4.10, 1.063, 0.298 and 0.670. Compared to PARIS, our method exhibits smoother surface generation results, as shown in Fig. 5. It is mainly because PARIS lacks a reasonable geometry initiation and directly utilizes the composite rendering, leading to the spatial overlap between the neural radiation fields of the start and end states. In contrast, our approach utilizes the accurate dense points from the reconstruction stage and optimizes the deformable fields to yield dynamic results, maintaining high-quality surface during generation.
It can be observed that our method is significantly superior to D-3DGS on most categories. As presented in Fig. 5, D-3DGS fails to accurately generate the surface of unseen states only using merely two states of RGB supervision. This is attributed to the fact that its deformation fields are exclusively derived from MLP. In contrast, our method accurately recovers the unseen states through the motion constraints.
4.4 Ablation Studies
@@
We also conduct an ablation study of the motion constraints on PartNet-Mobility dataset and present the quantitative results in Table. 4. For the baseline without motion constraints, we employ three MLPs to learn the variations of spatial positions, scales, and quaternions respectively for Gaussian primitives, which is similar to D-3DGS. The significantly superior results demonstrate the effectiveness of the motion constraints. Actually, motion constraints effectively compress the parameter search space of the deformation fields according to the kinematic structures of articulated objects and accurately infer the deformed spatial positions via rigid transformation.
@@
Table 4: The ablation study of motion constraints. We compare the surface generation quality on the PartNet-Mobility dataset between training with motion constraints and without motion constraints.
4.5 Generalization to the Real World
We conduct reconstruction and generation experiments in the real world to investigate the generalization capacity of our REArtGS. We report the quantitative results comparing with PARIS in Table. 5, and present the qualitative results in Fig. 6. Our approach significantly outperforms PARIS in mean performance across all metrics. This demonstrates that our method exhibits strong generalization capability for real-world objects and enables high-quality surface reconstruction and generation as shown in Fig. 6.
@@ 4.5 Generalization to the Real World
Box PARIS [15] 4.69 1.651 0.097 0.843 3.98 2.131 0.306 1.055
ArtGS [18] 8.61 0.967 0.143 0.697 10.23 1.606 0.309 0.851
@@
Stapler PARIS [15] 0.23 0.104 0.551 0.165 0.47 0.198 0.540 0.194
ArtGS [18] 51.04 1.364 0.476 0.827 5.13 0.527 0.533 0.514
@@
Scissor PARIS [15] 0.18 0.302 0.710 0.314 0.18 0.329 0.424 0.310
ArtGS [18] 11.43 0.073 0.588 0.192 5.53 0.089 0.600 0.213
@@
Cutter PARIS [15] 2.56 1.371 0.570 0.302 45.02 1.684 0.309 1.183
ArtGS [18] 5.65 0.453 0.644 0.477 65.64 1.298 0.093 0.806
@@
Drawer PARIS [15] 30.63 4.568 0.062 1.577 33.94 3.865 0.091 1.430
ArtGS [18] 19.41 7.478 0.086 1.935 12.60 1.624 0.094 0.911
@@
Eyeglasses PARIS [15] 0.16 1.582 0.637 0.212 0.18 1.550 0.609 1.284
ArtGS [18] 27.49 1.384 0.557 0.832 12.12 1.592 0.747 0.955
@@
Mean PARIS [15] 6.41 1.336 0.438 0.569 13.96 1.626 0.380 0.707
ArtGS [18] 20.61 1.953 0.416 0.826 18.54 1.123 0.396 0.708

### v4

@@ 4.1 Experimental Setting
Datasets. We conduct surface reconstruction experiments on synthetic dataset PartNet-Mobility xiang2020sapien and real-world dataset AKB-48 liu2022akb to evaluate the reconstruction quality of our method. Please see the Appendix for more details.
Metrics. Chamfer Distance (CD), F1-score and Earth Mover’s Distance (EMD) zhang2022deepemd are usually adopted as evaluation metrics for surface quality. However, these methods typically focus on the whole surface, endowing the unseen areas (such as the bottom of the object) with considerable weights during evaluation. This is unreasonable since we tend to be more concerned with the upper hemisphere regions of articulated objects. Therefore, we sample point clouds from the fragments between the rays cast by test cameras and the surface mesh to calculate EMD and CD, defined as CD (rs). We also evaluate the F1-score and CD for the whole surface, defined as CD (ws), to comprehensively evaluate the surface quality. Note that the CD values are multiplied by 1000 and the distance threshold of F1-score is set to 0.4. The calculation of these metrics can be referred to in our Appendix.
All the experiments are conducted on a single RTX 4090 GPU. We also provide detailed implementation of our approach and baselines in the Appendix.
@@ 4.2 Mesh Reconstruction Performance
We report the quantitative results of mesh reconstruction in Table. 1 and qualitative results are shown in Fig. 4. Note that we implement ArtGS without depth supervision for a fair comparison. As can be observed, our REArtGS outperforms state-of-the-art approaches in the mean across all metrics with 3.79, 1.236, 0.294 and 0.695 respectively. Compared to A-SDF mu2021sdf and Ditto jiang2022ditto that exploits 3D inputs for articulated object surface reconstruction, our method still achieves higher reconstruction quality in the majority of categories. In terms of PARIS, GOF and ArtGS, our approach exhibits significantly smoother and clearer surfaces as shown in Fig. 4, benefiting from our enhanced geometric constraints.
@@ 4.3 Mesh Generation Performance
We report the quantitative results of mesh generation in Table. 2 and present the qualitative results Fig. 5. Please refer to our Appendix for more qualitative results. Our REArtGS achieves the best mean results across all metrics with 4.10, 1.063, 0.298 and 0.670. Compared to PARIS, our method exhibits smoother surface generation results, as shown in Fig. 5. It is mainly because PARIS lacks a reasonable geometry initiation and directly utilizes the composite rendering, leading to the spatial overlap between the neural radiation fields of the start and end states. It can be also observed that our method is superior to ArtGS on most categories. This is primarily attributed to our high-quality geometry initialization from the reconstruction stage, as well as the motion-constrained deformable fields to yield dynamic generation.
Besides, we provide the part segmentation and joint parameter estimation results in the Appendix and  LABEL:sec:joint_parameter respectively.
@@
Table 4: The ablation study of motion constraints on PartNet-Mobility dataset.
@@
4.4 Ablation Studies
@@
Box PARIS liu2023paris 4.69 1.651 0.097 0.843 3.98 2.131 0.306 1.055
ArtGS liu2025building 8.61 0.967 0.143 0.697 10.23 1.606 0.309 0.851
@@
Stapler PARIS liu2023paris 0.23 0.104 0.551 0.165 0.47 0.198 0.540 0.194
ArtGS liu2025building 51.04 1.364 0.476 0.827 5.13 0.527 0.533 0.514
@@
Scissor PARIS liu2023paris 0.18 0.302 0.710 0.314 0.18 0.329 0.424 0.310
ArtGS liu2025building 11.43 0.073 0.588 0.192 5.53 0.089 0.600 0.213
@@
Cutter PARIS liu2023paris 2.56 1.371 0.570 0.302 45.02 1.684 0.309 1.183
ArtGS liu2025building 5.65 0.453 0.644 0.477 65.64 1.298 0.093 0.806
@@
Drawer PARIS liu2023paris 30.63 4.568 0.062 1.577 33.94 3.865 0.091 1.430
ArtGS liu2025building 19.41 7.478 0.086 1.935 12.60 1.624 0.094 0.911
@@
Eyeglasses PARIS liu2023paris 0.16 1.582 0.637 0.212 0.18 1.550 0.609 1.284
ArtGS liu2025building 27.49 1.384 0.557 0.832 12.12 1.592 0.747 0.955
@@
Mean PARIS liu2023paris 6.41 1.336 0.438 0.569 13.96 1.626 0.380 0.707
ArtGS liu2025building 20.61 1.953 0.416 0.826 18.54 1.123 0.396 0.708
@@
We also conduct an ablation study of the motion constraints on PartNet-Mobility dataset and present the quantitative results in Table. 4. For the baseline without motion constraints, we employ three MLPs to learn the variations of spatial positions, scales, and quaternions respectively for Gaussian primitives, which is similar to D-3DGS. The significantly superior results demonstrate the effectiveness of the motion constraints.
Figure 6: The qualitative results of surface reconstruction on real-world AKB-48 liu2022akb repository. We show both textured and non-textured meshes for two-part and multi-part articulated objects.
4.5 Generalization to the Real World
We conduct reconstruction and generation experiments in the real world to investigate the generalization capacity of our REArtGS. We report the quantitative results comparing with PARIS and ArtGS in Table. 5, and present the qualitative results in Fig. 6. Our approach significantly outperforms PARIS and ArtGS in mean performance across all metrics. This demonstrates that our method exhibits strong generalization capability for real-world objects and enables high-quality surface reconstruction and generation as shown in Fig. 6.
@@ 5 Conclusion and Future Work
Acknowledgements
This work was supported by National Natural Science Foundation of China under Grant 62302143, 62576130, and Anhui Province’s Key Science and Technology Project 202423k09020007.